OpenStudy (anonymous):

can somone help me slove this: y-x=-3 y+x=-3 by using elimination?

OpenStudy (angela210793):

add both parts 2y=-6 y=-3 -3+x=-3 x=-3+3=0 so the solution is (0;-3)